OpenStudy (anonymous):

Using the ratio of perfect squares method, what is square root of 59 rounded to the nearest hundredth? Do not use your calculator to answer this question. Use the ratio of perfect squares method.

OpenStudy (welshfella):

7^2 = 49 8^2 = 64

OpenStudy (anonymous):

can u explain?

OpenStudy (welshfella):

difference = 64-49 = 15 and we need square root of 59 59-49 = 10 so estimate of sqrt 59 = 7 + (10/15)

OpenStudy (welshfella):

we know the square root is between 7 and 8 and 59 is 10/15 of the difference between 49 and 64 so we use this fraction to add to the lower square root which is 7

OpenStudy (welshfella):

= 7.67

OpenStudy (welshfella):

which is quite close to value on the calculator (7.68)

OpenStudy (anonymous):

ohhh okay sorry im not good at math

OpenStudy (welshfella):

thats ok this is a learning website

OpenStudy (anonymous):

so 7.67 is the answer

OpenStudy (welshfella):

we pick out perfect squares either side of the number so if we required the square root of 29 we would use 5^2 = 25 and 6^2 = 36

OpenStudy (welshfella):

7,67 is the answer

OpenStudy (anonymous):

thanks :-)

OpenStudy (welshfella):

yw